OpenStudy (anonymous):

What formula could be correct for the speed "v" of ocean waves in terms of the density "p" of sea water, the acceleration of free fall "g", the depth of the ocean and the wavelength "y" ?

OpenStudy (anonymous):

Choices Please. I remember seeing this question somewhere. As far as my memory goes, I solved by dimensional analysis of the given options

OpenStudy (anonymous):

Sure thing.

OpenStudy (anonymous):

He he. Give me the complete question and I give you the sure answer

OpenStudy (anonymous):

with choices

OpenStudy (anonymous):

\[A) v=\sqrt{gy} \]

OpenStudy (anonymous):

\[ B) v=\sqrt{g/h} \]

OpenStudy (anonymous):

\[C) v= \sqrt{pgh}\]

OpenStudy (anonymous):

\[D) v=\sqrt{g/p}\]

OpenStudy (anonymous):

I would also appreciate an explanation :)

OpenStudy (anonymous):

Yes The option A is correct. Just remembered it now and yes I got it dimensional analysis

OpenStudy (anonymous):

How so?

OpenStudy (anonymous):

Wavelength(lamba) --> m g-->ms^(-2) \[\sqrt{gy} \] \[\sqrt{ms^{-2}m} \] \[\sqrt{m^2s^{-2}}\] \[ms^{-1}\] which is the unit of velocity. Try with rest of the options.It won't work :D

OpenStudy (anonymous):

It won't work means "It won't give dimensions as ms^(-1)"

OpenStudy (anonymous):

Ohhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhh That was so simple. Idk maybe I had a mindblock... too much physics ahhh!

OpenStudy (anonymous):

No problem. Make sure you eliminate options via Dimensional analysis if your brain is not working for a problem

OpenStudy (anonymous):

Right, thanks!

OpenStudy (anonymous):

Welcome :D